### v1

# Retrosynthesis Planning via Worst-path Policy Optimisation in Tree-structured MDPs

Abstract
Retrosynthesis planning aims to decompose target molecules into available building blocks, forming a synthesis tree where each internal node represents an intermediate compound and each leaf ideally corresponds to a purchasable reactant. However, this tree becomes invalid if any leaf node is not a valid building block, making the planning process vulnerable to the “weakest link” in the synthetic route. Existing methods often optimise for average performance across branches, failing to account for this worst-case sensitivity. In this paper, we reframe retrosynthesis as a worst-path optimisation problem within tree-structured Markov Decision Processes (MDPs). We prove that this formulation admits a unique optimal solution and offers monotonic improvement guarantees. Building on this insight, we introduce Interactive Retrosynthesis Planning (InterRetro), a method that interacts with the tree MDP, learns a value function for worst-path outcomes, and improves its policy through self-imitation, preferentially reinforcing past decisions with high estimated advantage. Empirically, InterRetro achieves state-of-the-art results, solving 100% of targets on the Retro*-190 benchmark, shortening synthetic routes by 4.9%, and achieving promising performance using only 10% of the training data - representing a significant advance in computational retrosynthesis planning.
1 Introduction
Retrosynthesis aims to identify the reactants needed to synthesise a target molecule with desired properties. As a fundamental task in computer-aided molecular design, retrosynthesis underpins progress in drug discovery and materials science [4, 39]. A key challenge in retrosynthesis is single-step prediction, which involves predicting the reactants for a given product (illustrated in Figure 1). While recent approaches using supervised learning have achieved human-level accuracy in this task [2, 38], their practical applicability is limited, as the suggested reactants are often commercially unavailable. Unlike single-step prediction, multi-step planning recursively decomposes the target into simpler intermediates, aiming to construct a synthetic route whose leaf nodes are purchasable compounds. This process naturally forms a sequential decision-making problem, where early decisions influence future steps and the overall outcome [7, 21].
Most methods tackle this decision-making problem using heuristic search, with Monte Carlo Tree Search (MCTS) being widely adopted for its ability to balance exploration and exploitation [17, 23]. In this framework, a single-step retrosynthesis model suggests candidate reactions and predicts the most likely reactants for a given product. During simulations, MCTS selects reactions by weighing their estimated value against their exploration potential. After simulating a complete synthetic route, the values of the decisions made along the path are updated to guide future searches more effectively. Building on this foundation, recent studies have proposed various enhancements to improve the exploration–exploitation trade-off or to stabilise value estimation [1, 7, 31, 37]. However, these methods heavily rely on time-consuming real-time search, requiring hundreds of model calls for each molecule, limiting their practical utility in large-scale molecular design scenarios.
To improve search efficiency, recent work fine-tunes pre-trained single-step models by imitating decisions made during successful heuristic search trajectories [9, 10, 34]. The key insight is that reaction choices selected after look-ahead planning are often more reliable than those proposed by the original model [22, 28]. By training on these improved decisions, the resulting policies can propose more plausible reactions and accelerate the search for viable synthetic routes. However, this fine-tuning process adapts the model to the distribution of molecules encountered during search, rather than those seen in direct inference. As a result, the model may perform well when used within the search loop but struggle to construct full synthetic routes independently — limiting its utility in settings where fast, search-free planning is desired.
Figure 1: Single-step prediction decomposes a molecule into reactants, whereas multi-step planning searches for a synthetic route, aiming to reach purchasable building blocks.
In this paper, we propose a novel perspective: reframing retrosynthesis planning as a worst-path optimisation problem in tree-structured Markov Decision Processes (tree MDPs). We observe that existing approaches typically optimise for average or cumulative rewards across all root-to-leaf paths in the synthesis tree [10, 21], overlooking a critical insight: a synthetic route is only valid if every root-to-leaf path terminates at a purchasable compound. Even a single unsuccessful path invalidates the entire synthetic route, making the worst-performing path the limiting factor in overall performance. This observation leads us to introduce a new “worst-path” objective that focuses explicitly on improving the most challenging path in the synthetic route.
Building on this novel formulation, we develop Interactive Retrosynthesis Planning (InterRetro) — a framework for multi-step retrosynthesis that learns to generate complete synthetic routes without search at inference time. InterRetro treats a single-step model as an agent operating within the tree MDP, recursively decomposing molecules into reactants through environment interactions. The agent is improved via weighted self-imitation of its past successful decisions to encourage shallow synthesis trees, effectively bootstrapping its performance. In particular, InterRetro operates in three key steps. Firstly, the agent interacts with the tree MDP to construct complete synthetic routes. Secondly, it identifies successful subtrees whose leaf nodes correspond to commercially available compounds. Finally, it fine-tunes the policy to imitate decisions from these subtrees, with theoretical guarantees of monotonic improvement and convergence to a unique optimal value function. This iterative self-improvement allows InterRetro to eliminate the need for real-time search while maintaining high planning quality.
This paper makes four key contributions: (1) we introduce the novel worst-path optimisation framework for tree-structured MDPs, specifically designed for problems where the weakest component determines the overall success; (2) we develop a weighted self-imitation learning algorithm with monotonic improvement guarantees for optimising this worst-path objective, proving the existence of a unique optimal solution through Bellman optimality analysis; (3) we apply this framework to retrosynthesis through InterRetro, a search-free approach to multi-step planning; and (4) we empirically demonstrate that InterRetro outperforms state-of-the-art (SOTA) methods in terms of success rate (achieving up to 100% on benchmark datasets), route length (reducing steps by 4.9%), and sample efficiency (reaching 92% of full performance with only 10% of training data).
2 Related Work
Single-step Prediction.
The fundamental task in retrosynthesis is single-step prediction, which identifies the reactants that produce a given target molecule. Early approaches are template-based, where reaction templates are extracted from patents and literature data [20], and a model selects the most appropriate template to decompose a molecule [2, 3]. Later methods propose template-free strategies, directly mapping the SMILES string of the target molecule to that of the reactants without relying on predefined transformation patterns [8, 26]. Recent studies introduce semi-template-based approaches, which combine both paradigms by first identifying intermediate structures and then completing them into full reactants — either by generating leaving groups [25], SMILES tokens [5, 32], or graph edits [30, 38]. While these single-step methods achieve high accuracy in predicting known reactions, they typically optimise for likelihood of historical reactions rather than synthesisability from commercially available compounds. In contrast, our approach explicitly favours reactions that lead to purchasable building blocks, enhancing downstream planning efficiency and practical applicability.
Multi-step Planning.
Multi-step planning aims to construct a synthesis tree rooted at the target molecule, expanded by a single-step model, and terminating at commercially available compounds. Existing methods rely on heuristic search to find viable synthetic routes. For example, [6] formulates the retrosynthesis problem as an AND/OR tree search and adopts proof-number search to find the optimal solution; Retro* [1] explores the AND/OR tree using A* search and proposes value estimation methods in this setting; To improve exploration, [23] and [37] adopt MCTS, which better balances exploitation and exploration compared to A*. These search-based approaches, however, require extensive computation at inference time, often necessitating hundreds of model calls per molecule. Recent efforts such as [9] and [10] address this limitation by fine-tuning the single-step model using successful trajectories collected during the planning phase, reducing search iterations and improving synthesis success rates. Yet, these methods still ultimately rely on search during inference, merely reducing rather than eliminating this computational burden.
Our method is closely related to [21], which frames retrosynthesis as a sequential decision process and learns a policy via RL. While prior methods typically optimise the cumulative reward across the entire synthesis tree and use value-based criteria to select reactions, we take a fundamentally different approach. We introduce a novel worst-path objective that focuses on the most challenging branch of the synthesis tree, and develop a weighted imitation learning algorithm that constrains the model’s support and promotes shallower synthesis trees. This formulation enables our method to generate complete synthetic routes without any search at inference time, while maintaining high success rates.
Self-imitation.
Self-imitation learning improves policy performance by encouraging the agent to replicate its own high-return past behaviours [15, 16]. A central aspect of self-imitation is the use of support constraints — the learnt policy is regularised to remain close to the data-generating policy, which stabilises training and mitigates distributional shift [16, 29]. This principle has been effectively applied in offline and offline-to-online reinforcement learning [14, 24, 27], where self-imitation facilitates safe policy improvement without extensive exploration. In our work, we apply self-imitation to retrosynthesis planning by initialising the single-step model to generate reactions from the dataset and refining it through its past successful decisions, ensuring that its proposed reactions remain chemically plausible throughout training while progressively favouring those that lead to commercially available building blocks.
3 Retrosynthesis via Worst-path Optimisation in Tree MDPs
Prior work in multi-step retrosynthesis typically minimises the total cost of a synthetic route by optimising cumulative rewards across all root-to-leaf paths in the synthesis tree. In contrast, our worst-path objective targets the weakest path, ensuring viability by requiring all paths to terminate at purchasable compounds. To efficiently optimise this criterion, we propose a weighted self-imitation algorithm that leverages successful trajectories while prioritizing improvement on failure-prone paths.
3.1 Tree-structured MDPs
We formalise retrosynthesis planning as a tree-structured Markov Decision Process (tree MDP), denoted by $\langle\mathcal{S},\mathcal{A},\mathcal{T},r,\mathcal{S}_{bb}\rangle$. Each state $s\in\mathcal{S}$ represents a molecule, each action $a\in\mathcal{A}$ represents a chemical reaction. Unlike standard MDPs where a transition yields a single successor, in retrosynthesis a reaction decomposes a molecule into multiple reactants. To capture this, we denote the set of all subsets of $\mathcal{S}$ (power set) as $2^{\mathcal{S}}$, and define a branching transition function $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow 2^{\mathcal{S}}$, which maps a molecule $s$ and a reaction $a$ to a finite set of reactants. These reactants become the children of $s$ in the synthesis tree. The reward function $r:\mathcal{S}\rightarrow\mathbb{R}$ assigns a numerical reward to each state, and $\mathcal{S}_{bb}\subset\mathcal{S}$ denotes the set of commercially available building blocks. A synthesis tree is successful only if all of its leaf nodes belong to $\mathcal{S}_{bb}$. A policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ specifies a probability distribution over actions for each state $s$, determining how the retrosynthesis tree expands. Throughout this paper, we denote a complete retrosynthesis tree by $\tau$ and let $P(\tau)$ represent the set of all root-to-leaf paths in $\tau$.
Figure 2: Distribution of reactant counts in the USPTO-50k dataset.
3.2 Worst-path Objective
The width of synthesis trees is naturally bounded as most molecules can be synthesised from a small number of reactants. As shown in Figure 2, $98.6\%$ of reactions in the USPTO-50k dataset involve three or fewer reactants. This observation indicates that route quality is primarily determined by depth rather than branching factor. Therefore, we evaluate a synthesis tree by its longest path from root to leaf. While our objective emphasises depth, it inherently addresses width by requiring that all reactants from a chosen reaction must lead to viable continuations. Figure 3 illustrates this concept: the left tree successfully synthesises molecule $\mathrm{A}$ with its value determined by the longest path, while the right tree fails due to an unsynthesisable intermediate, resulting in a return of 0.
Formally, we define a reward function, $r(s)=\mathds{1}\bigl{(}s\in\mathcal{S}_{bb}\bigr{)}$, which assigns a reward of $1$ if molecule $s$ is a commercially available building block, and 0 otherwise. Since $r(s_{t})$ is non-zero only at terminal states (i.e., leaf nodes), any root-to-leaf path $p=(s_{0},a_{0},s_{1},\ldots,s_{T})$ has return
$\sum_{t=0}^{T}\gamma^{t}\,r(s_{t})=\gamma^{T}\,r(s_{T}),$ (1)
where $\gamma\in(0,1)$ is the discount factor. Thus a successful path (with $r(s_{T})=1$) receives value $\gamma^{T}$, sharply penalising deeper decompositions. Dead-end paths (where $r(s_{T})=0$) have zero return. We then define our worst-path objective as
$J(\pi)=\mathbb{E}_{\tau\sim\pi}\left[\min_{p\in P(\tau)}\sum_{t=0}^{T}\gamma^{t}r(s_{t})\right].$ (2)
This objective captures the intuition that a synthetic route is only as viable as its weakest link: if any path leads to a dead-end, the entire tree’s value is zero. Maximising this worst-path return ensures every branch terminates at building blocks via the shortest possible routes. We also show the benefit of the worst-path objective over a best-path objective through an illustrative example in Appendix A.
Figure 3: Examples illustrating the tree MDP formulation. Each non-leaf node represents a molecule that is decomposed into one or more reactants. Left tree: A successful synthetic route for target molecule $\mathrm{A}$. It contains 44 root-to-leaf paths: $P(\tau)=\{\mathrm{ABD},\mathrm{ABEF},\mathrm{ABEGH},\mathrm{AC}\}$. Since all leaf nodes are building blocks, each path receives a value of $\gamma^{T}$ where $T$ is the path length. The tree’s overall value is $\min_{p\in P(\tau)}\{\gamma^{2},\gamma^{3},\gamma^{4},\gamma\}=\gamma^{4}$, determined by the longest path. Right tree: A failed synthesis attempt for molecule $\mathrm{A}$. One of its paths, $\mathrm{ABEG}$, terminates at $\mathrm{G}$, which is not a building block. This gives path $\mathrm{ABEG}$ a value of 0, making the tree’s overall value $\min_{p\in P(\tau)}\{\gamma^{2},\gamma^{3},0,\gamma\}=0$, illustrating why a single failing path invalidates the entire route.
3.3 Value Estimation
To optimise the worst-path objective, the agent interacts with the tree MDP and imitates its advantageous past decisions. We quantify the advantage of a decision through a Q-value function $Q^{\pi}(s,a)$: given a molecule $s$ as the root node, it estimates the expected worst-path return when first expanding the molecule with reaction $a$, and thereafter following policy $\pi$. Formally:
$Q^{\pi}(s,a)=\mathbb{E}_{\tau\sim\pi}\left[\min_{p\in P(\tau)}\sum_{t=0}^{T}\gamma^{t}r(s_{t})\;\Big{|}\;s_{0}=s,a_{0}=a\right],$ (3)
where $s_{0}$ and $a_{0}$ denote the root node and the reaction applied to it, respectively. Similarly, the value function $V^{\pi}(s)$ represents the expected worst-path return when expanding a molecule $s$ using policy $\pi$ throughout:
$V^{\pi}(s)=\mathbb{E}_{\tau\sim\pi}\left[\min_{p\in P(\tau)}\sum_{t=0}^{T}\gamma^{t}r(s_{t})\;\Big{|}\;s_{0}=s\right].$ (4)
The advantage function quantifies the relative benefit of applying reaction $a$ to molecule $s$ compared to following policy $\pi$:
$A^{\pi}(s,a)=Q^{\pi}(s,a)-V^{\pi}(s).$ (5)
A positive advantage indicates that reaction $a$ leads to better outcomes than the policy’s average behaviour. The key insight of our method is to prioritise imitation of past decisions with high advantages. The following proposition establishes a relationship between Q-value and value functions, enabling more efficient advantage estimation:
Proposition 1.
The Q-value function $Q^{\pi}(s,a)$ equals its immediate reward plus the discounted value of its next states:
$Q^{\pi}(s,a)=r(s)+\gamma(1-r(s))\min_{s^{\prime}\in\mathcal{T}(s,a)}V^{\pi}(s^{\prime})$ (6)
The proof is provided in Appendix C.1.
Based on this proposition, we can estimate the advantage as:
$A^{\pi}(s,a)=r(s)+\gamma(1-r(s))\min_{s^{\prime}\in\mathcal{T}(s,a)}V^{\pi}(s^{\prime})-V^{\pi}(s).$ (7)
Furthermore, Proposition 2 shows that $V^{\pi}(s)$ satisfies a recursive relation, providing the basis for value function approximation:
Proposition 2.
The value function $V^{\pi}(s)$ satisfies the recursion:
$V^{\pi}(s)=r(s)+\gamma(1-r(s))\sum_{a\in\mathcal{A}}\pi(a\mid s)\min_{s^{\prime}\in\mathcal{T}(s,a)}V^{\pi}(s^{\prime}).$ (8)
The proof is provided in Appendix C.2.
Beyond the value functions for a given policy, we can define the optimal worst-path value function, $V^{*}$, which represents the maximum possible worst-path return achievable from any state.
Proposition 3.
This optimal value function $V^{*}$ uniquely satisfies the Bellman optimality equation for the worst-path objective:
$V^{*}(s)=r(s)+\gamma(1-r(s))\max_{a\in\mathcal{A}}\left[\min_{s^{\prime}\in\mathcal{T}(s,a)}V^{*}(s^{\prime})\right],$ (9)
The proof is provided in Appendix C.3.
In the proof, we show that the corresponding Bellman optimality operator is a contraction mapping, which guarantees the existence and uniqueness of $V^{*}$ and the convergence of value iteration to $V^{*}$. Furthermore, there exists at least one deterministic stationary policy $\pi^{*}$ that is greedy with respect to $V^{*}$ and is therefore an optimal policy.
3.4 Self-imitation Learning for Retrosynthesis
A key challenge in retrosynthesis is ensuring that the learnt policy proposes realistic chemical reactions at each step. To accomplish this, we leverage a pretrained single-step model $\pi^{0}$ that reflects expert or empirical knowledge from the training set. Single-step models empirically produce valid chemical transformations with high fidelity, motivating us to maintain their support when learning an improved policy. Specifically, we define the constrained policy set:
$\Pi=\{\pi\mid\pi(a\mid s)=0\text{ whenever }\pi^{0}(a\mid s)=0\}.$ (10)
By restricting $\pi$ to actions that $\pi^{0}$ assigns non-zero probability, we eliminate the risk of generating unrealistic reactions while allowing flexibility to re-weight feasible actions based on their effectiveness for multi-step planning.
Our objective is to find a policy $\pi\in\Pi$ that maximises the worst-path objective. We achieve this through an iterative procedure. In each iteration $i$, we aim to find an improved policy $\pi^{i+1}$ by imitating advantageous state-action pairs $(s,a)$ experienced under policy $\pi^{i}$. The advantage $A^{\pi^{i}}(s,a)$ quantifies this, and the learning objective for $\pi^{i+1}$ is formulated as:
$J(\pi^{i+1})=\mathbb{E}_{s\sim d_{\pi^{i}}(\cdot),a\sim\pi^{i}(\cdot\mid s)}\left[\exp\bigl{(}\beta A^{\pi^{i}}(s,a)\bigr{)}\log\pi^{i+1}(a\mid s)\right],$ (11)
where $\beta>0$ is the advantage coefficient controlling the strength of advantage weighting, and $d_{\pi^{i}}$ is the state distribution induced by policy $\pi^{i}$ [29]. Intuitively, reactions with higher advantages receive higher weights, guiding the policy toward better-than-average reactions.
Because each new policy $\pi^{i+1}$ is derived by re-weighting $\pi^{i}$, and the initial policy $\pi^{0}$ restricts the support, the entire policy sequence $\{\pi^{i}\}_{i\geq 0}$ remains within the feasible set $\Pi$ [12]. This ensures that all proposed reactions throughout training remain chemically valid.
We now show that this self-imitation update leads to monotonic improvement in the worst-path value function:
Proposition 4.
Let $\pi^{i+1}$ be the policy obtained by optimising the objective in Eq.11. Then, the updated policy is guaranteed to perform at least as well as the previous policy for all states:
$V^{\pi^{i+1}}(s)\geq V^{\pi^{i}}(s),\forall s\in\mathcal{S}.$ (12)
The proof is provided in Appendix C.4.
Proposition 4 guarantees monotonic improvement under the exact policy update:
$\pi^{i+1}(a\mid s)\propto\pi^{i}(a\mid s)\exp\bigl{(}\beta A^{\pi^{i}}(s,a)\bigr{)},$ (13)
This update re-weights the previous policy by exponentiated advantages, increasing the likelihood of actions that outperform the current expectation.
For policy optimisation, we use the current policy $\pi^{i}$ to interact with the retrosynthesis environment and collect data to learn the next iteration’s policy $\pi^{i+1}$. Through iterative weighted imitation, the policy increasingly favours high-quality reactions that lead to successful synthetic routes with shorter paths, while maintaining chemical plausibility by respecting the support of the pretrained model.
4 The InterRetro Algorithm
Algorithm 1 Interactive retrosynthesis planning (InterRetro).
0: Pretrained one-step policy $\pi_{\theta}$, value function $V_{\phi}$, training dataset $\mathcal{D}$, replay buffer $\mathcal{B}$.
0:Explore($\pi_{\theta}$, $m$):
[…]
4:while $\texttt{q}\neq\emptyset$ and $\texttt{step}<\texttt{max\_steps}$ do
5:  $s\leftarrow\texttt{q.pop()}$
6:  $a,\mathcal{S}_{r}\leftarrow\pi_{\theta}.\texttt{get\_reactants}(s)$
7:  $\texttt{tree.add\_branch}(s,a,\mathcal{S}_{r})$
8:
9:  # Add non-building blocks
10:  $\texttt{q}\leftarrow\texttt{q}\cup\{\,s^{\prime}\in\mathcal{S}_{r}\mid s^{\prime}\notin\mathcal{S}_{bb}\,\}$
11:
12:  $\texttt{step}\leftarrow\texttt{step}+1$
13:end while
14:return  tree
0:InterRetro($\pi_{\theta}$, $m$):
1:for $i=1,\dots,I$ do
2:  while $\mathcal{D}$ is not empty do
3:   $m\leftarrow\mathcal{D}.\texttt{pop()}$
4:   $\texttt{tree}\leftarrow\textsc{Explore}(\pi_{\theta},m)$
5:   $\texttt{brs}\leftarrow\{\}$
6:   for each subtree $\tau\in\texttt{tree}$ do
7:    if $\tau$ is successful then
8:     $\texttt{brs}\leftarrow\texttt{brs}\cup\textsc{AllBranches}(\tau)$
9:    end if
10:   end for
11:   $\mathcal{B}.\texttt{append}(\texttt{brs})$
12:   $\texttt{branches}\leftarrow\mathcal{B}.\texttt{sample()}$
13:   $V_{\phi}.\texttt{update}(\texttt{branches})$ $\triangleright$ Eq. 14
14:   $\pi_{\theta}.\texttt{update}(V_{\phi},\ \texttt{branches})$ $\triangleright$ Eq. 15
15:  end while
16:end for
We now present our algorithm InterRetro for retrosynthesis planning. Section 4.1 introduces how the single-step model interacts with the tree MDP and collects experiences. Section 4.2 describes how to learn a value function on the worst-path objective and how to fine-tune the policy to reproduce advantageous decisions. Finally, Section 4.3 provides more implementation details for reproducibility.
4.1 Environment Interactions
The Explore procedure in Algorithm 1 shows how InterRetro constructs a synthesis tree by interacting with the tree MDP. Starting from a target molecule $m$, the single-step model proposes a reaction $a$ to decompose the molecule into a set of reactants $\mathcal{S}_{r}$. These reactants will be attached to the parent node $m$ in the synthesis tree. Among them, the non-building blocks are then placed in a collection (e.g., a queue) for further expansion. Each subsequent round pops a molecule from the collection to continue the decomposition process. The interaction terminates when the collection becomes empty (meaning all leaf nodes are building blocks) or when a predefined maximum number of steps is reached.
4.2 Value Function and Policy Learning
The Explore procedure constructs synthesis trees by interacting with the tree MDP. We extract all branches $(s,a,\mathcal{S}_{r})$ within successful subtrees and store them in a first-in first-out (FIFO) replay buffer $\mathcal{B}$ for value function and policy learning (see InterRetro in Algorithm 1).
To learn the value function network $V_{\phi}$ with parameter $\phi$, we minimise the mean squared error between the predicted value $V_{\phi}(s)$ and its Bellman target derived from Proposition 2:
$\mathcal{L}(\phi)=\mathbb{E}_{(s,a,\mathcal{S}_{r})\sim\mathcal{B}}\left[(V_{\phi}(s)-(r(s)+\gamma(1-r(s))\min_{s^{\prime}\in\mathcal{S}_{r}}V_{\phi^{-}}(s^{\prime})))^{2}\right],$ (14)
where $V_{\phi^{-}}$ is a target network updated slowly for stability.
The policy network $\pi_{\theta}$ with parameter $\theta$ is updated using the weighted imitation learning objective from Eq. 11, implemented as the following loss function:
$\mathcal{L}(\theta)=\mathbb{E}_{(s,a,\mathcal{S}_{r})\sim\mathcal{B}}\left[\exp_{\mathrm{clip}}(\beta A_{\phi}(s,a))\log\pi_{\theta}(a\mid s)\right],$ (15)
where $\beta>0$ is the advantage coefficient which controls the imitation strength on the past successful experience, and $\exp_{\mathrm{clip}}(\cdot)$ is the exponential function with a clipped output range $(0,C]$ for numerical stability. The advantage is estimated by the value function network, according to Proposition 1:
$A_{\phi}(s,a)=r(s)+\gamma(1-r(s))\min_{s^{\prime}\in\mathcal{T}(s,a)}V_{\phi}(s^{\prime})-V_{\phi}(s).$ (16)
This joint optimisation of the value and policy networks enables InterRetro to iteratively improve its planning policy while maintaining robustness and chemical validity.
4.3 Implementation Details
We choose Graph2Edits as the single-step model due to its strong performance and flexibility in single-step retrosynthesis prediction [38]. Graph2Edits represents a molecule through a graph and predicts a sequence of graph edits to transform it to reactants. These graph edits can delete bonds, modify bond types, alter atoms, or attach leaving groups (see Appendix D for more details). The single-step model and our proposed value function encode the molecule using a message passing neural network [33] and forward the latent code into linear layers with ReLU as the activation function. For training, we run $6$ parallel exploration processes and collect $36$ synthetic trees per iteration. The networks are updated $5$ times per round using data from a compact replay buffer of $20,000$ branches, chosen to reduce CPU memory usage and maintain close alignment between the data-collecting policy $\pi^{i}$ and the updated policy $\pi^{i+1}$. Our models are trained on a single NVIDIA RTX A5000 GPU and, without pre-training the single-step model, require approximately $48$ hours to fully converge. The code will be open-sourced after the publication of this manuscript.
5 Experimental Results
In this section, we aim to answer the following questions: (1) What are the advantages of our proposed method compared to the SOTA algorithms? (2) How does each component of the method contribute to the performance? Additionally, we examine the real-world feasibility of the proposed synthetic routes and present illustrative examples in Appendix B.
5.1 Benchmark Results
The proposed method is trained by creating synthetic routes for nearly $300k$ molecules in the USPTO-50k dataset with commercially available building blocks from the eMolecules dataset111eMolecules: https://downloads.emolecules.com/free/.. We evaluate performance on three benchmarks of increasing difficulty: Retro*-190 [1], ChEMBL-1000, and GDB17-1000 [10], where the suffix indicates the dataset size. ChEMBL-1000 from ChEMBL dataset [35] is a manually curated collection of bioactive molecules exhibiting drug-like properties, while GDB17-1000 is a subset of the GDB-17 dataset [18], which covers a size range common for typical drugs and lead compounds. We compare against established retrosynthesis planning methods: MCTS, Retro* [1], Self-improve [9], PDVN [10], and GraphRetro [31]. Additionally, we include two single-step methods, MEGAN [19] and Graph2Edits [38], combined with Retro* as search algorithm, and two recently proposed methods, DreamRetroer [36] and RetroCaptioner [11]. Baseline results are produced from their official implementations or the Syntheseus project [13].
Success Rate.
Success rate measures the percentage of target molecules that can be successfully decomposed into building blocks. In Table 1, we firstly compare methods under different model-call budgets: $100$, $200$, and $500$. Our proposed InterRetro significantly outperforms SOTA algorithms across all three test sets. With $500$ model calls, InterRetro achieves $100\%$, $98.2\%$, and $99.5\%$ success rates on Retro*-190, ChEMBL-1000, and GDB17-1000, respectively. All synthetic routes generated by InterRetro on Retro*-190 using 500 model calls are provided in the supplementary materials.
Table 1: Performance evaluation on three benchmarks. The evaluation metrics include the success rate under different test molecules with different budget of model calls, which are direct generation (DG), $100$, $200$ and $500$ model calls. The DG columns are single-step model’s results without search.
Retro-190 ChEMBL-1000 GDB17-1000 Single-step Search DG 100 200 500 DG 100 200 500 DG 100 200 500 Template MCTS $20.00$ $43.68$ $47.37$ $62.63$ $32.00$ $45.60$ $68.80$ $71.90$ $3.00$ $3.20$ $3.70$ $4.50$ Template Retro* $20.00$ $38.42$ $58.42$ $79.47$ $32.00$ $69.10$ $72.00$ $74.70$ $3.00$ $5.40$ $6.60$ $7.50$ LocalRetro MCTS 22.10 $44.21$ $57.36$ $62.10$ $47.30$ $62.70$ $69.10$ $75.00$ $4.60$ $14.00$ $16.70$ $20.30$ LocalRetro Retro* $22.10$ $58.94$ $64.73$ $73.68$ $47.30$ $74.80$ $80.40$ $82.40$ $4.60$ $18.90$ $22.20$ $28.80$ MEGAN Retro* $8.42$ $60.52$ $62.10$ $73.15$ $38.00$ $71.70$ $75.40$ $79.00$ $6.20$ $37.60$ $45.70$ $57.20$ Graph2Edits Retro* $16.84$ $41.05$ $50.00$ $56.31$ $47.10$ $68.70$ $78.80$ $80.70$ $5.90$ $18.20$ $24.00$ $32.20$ Self-improve Retro* - $67.37$ $83.16$ $96.32$ - - - $81.10$ - - - $15.00$ PDVN Retro* - $93.68$ $97.37$ $98.95$ - - - $83.50$ - - - $26.90$ RetroCaptioner Retro* $5.26$ $68.94$ $72.63$ $85.26$ $3.90$ $72.60$ $76.50$ $78.70$ $3.20$ $56.20$ $68.20$ $75.20$ DreamRetroer $32.10$ $78.94$ $88.42$ $90.52$ $31.10$ $78.10$ $81.70$ $83.10$ $4.20$ $27.36$ $28.97$ $33.20$ InterRetro MCTS $\boldsymbol{95.78}$ $89.47$ $98.94$ $\boldsymbol{100.00}$ $\boldsymbol{93.10}$ $78.40$ $89.30$ $97.50$ $\boldsymbol{89.00}$ $80.80$ $96.10$ 99.50 InterRetro Retro* $\boldsymbol{95.78}$ $\boldsymbol{96.31}$ $\boldsymbol{100.00}$ $\boldsymbol{100.00}$ $\boldsymbol{93.10}$ $\boldsymbol{91.40}$ $\boldsymbol{96.20}$ $\boldsymbol{98.20}$ $\boldsymbol{89.00}$ $\boldsymbol{83.80}$ $\boldsymbol{96.50}$ $97.20$
Most notably, InterRetro maintains exceptional performance on the challenging GDB17-1000 benchmark, where prior approaches such as PDVN and Self-improve struggle ($26.9\%$ and $15.0\%$ respectively). This dramatic performance gap suggests that our worst-path objective effectively handles complex molecules that require precise reaction planning.
Furthermore, InterRetro can directly generate high-quality synthetic routes without any search algorithm, as shown in the Direct Generation (DG) columns. Our search-free performance ($95.78\%$, $93.10\%$, and $89.00\%$ across the three benchmarks) substantially exceeds even the search-based performance of competing methods. This demonstrates that our self-imitation learning approach successfully transfers planning capabilities to inference time, effectively eliminating the computational bottleneck of real-time search.
Table 2: The average length of the routes on the Retro*-190 test set.
Algorithm Average Length
Retro* 5.83
RetroGraph 5.63
PDVN 4.83
MEGAN 4.12
Graph2Edits 4.38
InterRetro 3.92
Route Length.
Route length is the number of reactions needed to synthesise a target molecule. Route length could be a concern with our worst-path objective since it does not consider the width of the tree and the total number of reactions required. In Table 2, we compare the route length with other baselines on the $138$ target molecules that all methods can resolve. Our method outperforms the SOTA by $4.85\%$.
Figure 4: Experimental figures. (a) Performance under different training data usage and computation budgets. (b) Statistics on estimated depth of synthesis trees. (c) Ablations on the advantage coefficient.
Sample Efficiency.
Sample efficiency refers to the amount of training data required to learn an effective policy. We evaluate the policy using subsets of the training data at $1\%$, $2\%$, $5\%$, $10\%$, and $100\%$, corresponding to approximately $3k$, $6k$, $15k$, $30k$, and $300k$ molecules. As shown in Figure 4(a), InterRetro achieves near SOTA performance ($92\%$) with just $10\%$ of the training set when directly generating synthesis trees, and reaches SOTA performance when combined with $200$ Retro* search iterations. Notably, when trained on the full dataset, InterRetro outperforms most existing methods in direct generation, achieving a success rate of $95.78\%$. Similar trends are observed on the ChEMBL-1000 and GDB17-1000 benchmarks.
5.2 Ablation Studies
Value Estimation.
We proposed a value function to estimate the worst-path return. The worst-path return indicates the estimated depth of the synthesis tree: $\text{depth}(s)=\frac{\log V_{\phi}(s)}{\log\gamma}$. In Figure 4(b), we investigate the distribution of the estimated depth, corresponding to the real depth by direct generation. We found that the value function can reflect the difficulties to synthesise the molecules, while it shows better capability on the molecules that require a less deep synthetic route.
Advantage Coefficient.
The hyperparameter $\beta$ controls the imitation strength on the past successful experience. A large $\beta$ means a focused imitation on the action with a high advantage. Figure 4(c) shows the learning curve on the Retro*-190 test set with $\beta\in\{0,1,2,5,10,20\}$. The figure shows that the DG performance increases from $16\%$ to around $80\%$ by uniformly imitating successful past experiences. With the advantage weighting, the performance soars to a success rate of more than $90\%$ when $\beta$ increases to $2$, and more than $95\%$ with $\beta\geq 10$.
6 Discussion and Conclusions
We introduced InterRetro, a novel approach that frames retrosynthesis planning as worst-path optimisation in tree-structured MDPs. Our weighted self-imitation algorithm enables single-step models to generate high-quality synthetic routes without search at inference time, achieving SOTA performance in success rate, route length, and sample efficiency. Whilst this worst-path policy optimisation has proven effective, future investigations could explore alternative worst-path criteria, such as minimising the maximum depth among all successful paths, incorporating the cost of the reaction and the building blocks, or designing a heuristic search to optimise the worst-path objective.
Whilst our formulation offers significant advantages, it assumes that all proposed reactions are feasible in real-world settings — an assumption that does not always hold. Although we mitigate this by imitating only actions supported by pretrained models (similar to [1, 10]), computational approaches cannot guarantee practical feasibility without experimental validation. Real chemical reactions depend on specific conditions that are rarely captured in retrosynthesis datasets. Future work should therefore focus on creating related datasets and integrating reaction feasibility checks into InterRetro. Moreover, this “weakest link” principle is broadly applicable beyond chemistry to sequential decision problems where overall success depends on the weakest component, such as robust project planning where delays in any critical task affect the entire timeline, and multi-agent systems where team performance is limited by the least capable agent.
Appendix A Illustrative Example: Worst-Path vs. Best-Path Optimisation
To further clarify the motivation behind our worst-path objective, we present a simplified pedagogical example that illustrates the critical “weakest link” principle in retrosynthesis planning. Using the same tree-structured MDP framework (Section 3.1), we demonstrate why optimising for the worst-performing path is essential for practical synthetic route design.
Consider a scenario with discount factor $\gamma=0.9$ and a target molecule $M_{0}$ that can be addressed by two distinct initial reactions, $a_{1}$ and $a_{2}$. Each reaction produces reactants that must all be successfully processed for the overall route to be viable. We define two action-value functions for a state $s$ and action $a$:
$J_{WP}(s,a)$ $=\min_{p\in P(\mathcal{T}(s,a))}V(p),\quad$ (Worst-Path Action-Value) (17)
$J_{BP}(s,a)$ $=\max_{p\in P(\mathcal{T}(s,a))}V(p),\quad$ (Best-Path Action-Value) (18)
where $P(\mathcal{T}(s,a))$ is the set of all root-to-leaf paths in the tree generated by taking action $a$ from state $s$, and $V(p)$ is the discounted return of a specific path $p$. The Worst-Path (WP) objective evaluates actions by their worst-case outcome, whilst the Best-Path (BP) objective considers only the most favourable outcome.
$M_{0}$ $R_{1A}$ $R_{1B}$ $BB_{1}$ $E$ $a_{1}$ Tree $T_{1}$: $J_{WP}(M_{0},a_{1})=0$ $J_{BP}(M_{0},a_{1})=0.81$ $M_{0}$ $R_{2A}$ $R_{2B}$ $R^{\prime}_{2A}$ $R^{\prime}_{2B}$ $B_{2} BB$$B_{3}$ $a_{2}$ Tree $T_{2}$: $J_{WP}(M_{0},a_{2})=0.729$ $J_{BP}(M_{0},a_{2})=0.729$
Figure 5: An illustrative Tree MDP comparing two potential reaction pathways from target molecule $M_{0}$ (blue circles). Each reaction (action $a_{1}$ or $a_{2}$) decomposes $M_{0}$ into multiple reactants that must all be synthesised. Yellow circles labelled $BB_{i}$ represent purchasable building blocks that yield reward 1, while the red circle labelled $E$ represents a dead-end with reward 0. Tree $T_{1}$ (left) has one successful 2-step path and one failed path. Tree $T_{2}$ (right) has two successful 3-step paths. Path values $V(p)$ are discounted by $\gamma=0.9$, giving $0.9^{2}=0.81$ for 2-step paths and $0.9^{3}=0.729$ for 3-step paths.
Let $V(p)$ be the discounted return of a complete path $p$ starting from $M_{0}$ and ending at a leaf node, calculated as $\gamma^{L}r(\text{leaf})$, where $L$ is the total path length from $M_{0}$, and $r(\text{leaf})$ is 1 for a $BB$ and 0 for a dead-end $E$.
Evaluating Action $a_{1}$ (leading to Tree $T_{1}$):
  • Path $p_{1A}$ (via $R_{1A}$ to $BB_{1}$): Length $L=2$. $V(p_{1A})=\gamma^{2}\cdot 1=(0.9)^{2}=0.81$.
  • Path $p_{1B}$ (via $R_{1B}$ to $E$): Length $L=2$. $V(p_{1B})=\gamma^{2}\cdot 0=0$.
The worst-path action-value for $a_{1}$ is:
$J_{WP}(M_{0},a_{1})=\min\{V(p_{1A}),V(p_{1B})\}=\min\{0.81,0\}=0.$
Evaluating Action $a_{2}$ (leading to Tree $T_{2}$):
  • Path $p_{2A}$ (via $R_{2A}$, $R^{\prime}_{2A}$ to $BB_{2}$): Length $L=3$. $V(p_{2A})=\gamma^{3}\cdot 1=(0.9)^{3}=0.729$.
  • Path $p_{2B}$ (via $R_{2B}$, $R^{\prime}_{2B}$ to $BB_{3}$): Length $L=3$. $V(p_{2B})=\gamma^{3}\cdot 1=(0.9)^{3}=0.729$.
The worst-path action-value for $a_{2}$ is:
$J_{WP}(M_{0},a_{2})=\min\{V(p_{2A}),V(p_{2B})\}=\min\{0.729,0.729\}=0.729.$
Therefore, a policy $\pi_{WP}$ optimising the worst-path objective (Eq. 17) would choose action $a_{2}$ (value $0.729$) over action $a_{1}$ (value 0).
Now, considering the Best-Path Seeking objective (Eq. 18):
  • For action $a_{1}$: $J_{BP}(M_{0},a_{1})=\max\{V(p_{1A}),V(p_{1B})\}=\max\{0.81,0\}=0.81$.
  • For action $a_{2}$: $J_{BP}(M_{0},a_{2})=\max\{V(p_{2A}),V(p_{2B})\}=\max\{0.729,0.729\}=0.729$.
A policy $\pi_{BP}$ optimising this best-path objective would choose action $a_{1}$ (value $0.81$) over action $a_{2}$ (value $0.729$).
This example highlights a key difference: the Best-Path policy $\pi_{BP}$ is drawn to the higher potential reward of path $p_{1A}$ in $T_{1}$, despite the significant risk of complete failure associated with path $p_{1B}$. In contrast, our worst-path policy $\pi_{WP}$ strictly prefers the entirely viable option $a_{2}$ (albeit with individually longer paths), ensuring that all branches of the synthesis lead to success. This inherent risk aversion is crucial for practical retrosynthesis planning, where a single unresolvable intermediate invalidates the entire proposed route.
Appendix B Case Studies
Figure 6: Predicted synthetic routes on three randomly selected molecules in Retro*-190. The yellow circles highlight the reaction centres.
In Figure 6, we randomly select three targets from the Retro*-190 benchmark and illustrate their complete synthesis trees as directly generated by InterRetro in a single forward pass. (i) For the sulfonylated chiral cyclohexanol derivative, InterRetro proposes a four-step route involving: a sulfonylation of a chiral alcohol to introduce the sulfonate ester, an electrophilic bromination of the diazine core, a C(sp3)–C(sp2) cross-coupling to build the fused triazolopyrimidine system, and an N-arylation to complete the final structure. (ii) For the alkyne-substituted diazine, InterRetro constructs the target via two steps: first an acylation of the brominated azine to give the N-formyl intermediate, followed by an SN2 N-propargylation. (iii) For the complex triazolopyridine-fused drug-like scaffold, InterRetro suggests a three-step route: a benzylic etherification that assembles the protected piperidine core, a C(sp³)–C(sp²) cross-coupling to append the triazolopyridine fragment, and an N-alkylation that installs the silyl-protected side chain. Across all three cases, InterRetro applies feasible chemical transitions and terminates in commercially available building blocks. A full list of the synthetic routes on Retro*-190 is attached in the supplementary materials.
Appendix C Missing Proofs and Further Theory
C.1 Proof of Proposition C.1
Proof.
Recall that
$Q^{\pi}(s,a)\;=\;\mathbb{E}_{\tau\sim\pi}\!\Bigl{[}\min_{p\in P(\tau)}\sum_{t=0}^{T}\gamma^{t}\,r\bigl{(}s_{t}\bigr{)}\;\Big{|}\;s_{0}=s,\;a_{0}=a\Bigr{]}.$
Since the transition function $\mathcal{T}$ is deterministic, taking action $a$ in $s$ immediately leads to the set of next states $\mathcal{T}(s,a)$. Observe that $r(s)$ is $1$ if and only if $s$ is a building block (in which case no further transitions occur), and 0 otherwise. We prove the desired equality by a simple case distinction:
1. Case $\boldsymbol{r(s)=1}$ (i.e., $s$ is a building block). In this case, once we reach $s$, the path terminates and collects reward $1$. Hence
$Q^{\pi}(s,a)\;=\;\underbrace{r(s)}_{=1}\;=\;r(s)+\gamma\,\bigl{(}1-r(s)\bigr{)}\,\min_{s^{\prime}\in\mathcal{T}(s,a)}\,V^{\pi}(s^{\prime})\quad\text{(since $1-r(s)=0$).}$
2. Case $\boldsymbol{r(s)=0}$ (i.e., $s$ is not a building block). Any path extending from $(s,a)$ must proceed into one of the next states $s^{\prime}\in\mathcal{T}(s,a)$. Because we are taking a “worst-path” (minimum-return) perspective, the worst-case continuation value from $s$ under action $a$ is
$\min_{s^{\prime}\in\mathcal{T}(s,a)}V^{\pi}(s^{\prime}),$
and each step is discounted by $\gamma$. Therefore,
$Q^{\pi}(s,a)\;=\;\underbrace{r(s)}_{=0}\;+\;\gamma\,\min_{s^{\prime}\in\mathcal{T}(s,a)}\,V^{\pi}(s^{\prime})\;=\;r(s)\;+\;\gamma\,\bigl{(}1-r(s)\bigr{)}\,\min_{s^{\prime}\in\mathcal{T}(s,a)}\,V^{\pi}(s^{\prime}).$
Combining both cases completes the proof:
$Q^{\pi}(s,a)\;=\;r(s)\;+\;\gamma\,\bigl{(}1-r(s)\bigr{)}\,\min_{s^{\prime}\in\mathcal{T}(s,a)}V^{\pi}(s^{\prime}).$
C.2 Proof of Proposition 2
Proof.
Recall that
$V^{\pi}(s)\;=\;\mathbb{E}_{\tau\sim\pi}\!\Bigl{[}\min_{p\in P(\tau)}\sum_{t=0}^{T}\gamma^{t}\,r\bigl{(}s_{t}\bigr{)}\;\Big{|}\;s_{0}=s\Bigr{]}.$
When we are in state $s$, the next action $a$ is sampled from the policy $\pi(\cdot\mid s)$. Because the environment is deterministic, taking $(s,a)$ leads to the set of next states $\mathcal{T}(s,a)$. By the definition of the worst-path criterion (the inner minimum), we have
$Q^{\pi}(s,a)\;=\;r(s)\;+\;\gamma\,(1-r(s))\;\min_{s^{\prime}\in\mathcal{T}(s,a)}\,V^{\pi}(s^{\prime}),$
and
$V^{\pi}(s)\;=\;\sum_{a\in\mathcal{A}}\,\pi(a\mid s)\;Q^{\pi}(s,a).$
Substitute $Q^{\pi}(s,a)$ into the sum:
$V^{\pi}(s)\;=\;\sum_{a\in\mathcal{A}}\pi(a\mid s)\Bigl{[}r(s)\;+\;\gamma\,(1-r(s))\;\min_{s^{\prime}\in\mathcal{T}(s,a)}V^{\pi}(s^{\prime})\Bigr{]}.$
Since $\sum_{a\in\mathcal{A}}\pi(a\mid s)=1$, we can factor out $r(s)$ to obtain
$V^{\pi}(s)\;=\;r(s)\;\underbrace{\sum_{a\in\mathcal{A}}\pi(a\mid s)}_{=1}\;+\;\gamma\,(1-r(s))\,\sum_{a\in\mathcal{A}}\pi(a\mid s)\,\min_{\,s^{\prime}\in\mathcal{T}(s,a)}V^{\pi}(s^{\prime}).$
Hence,
$V^{\pi}(s)\;=\;r(s)\;+\;\gamma\,\bigl{(}1-r(s)\bigr{)}\sum_{a\in\mathcal{A}}\pi(a\mid s)\,\min_{\,s^{\prime}\in\mathcal{T}(s,a)}V^{\pi}(s^{\prime}),$
completing the proof. ∎
C.3 Bellman Optimality Operator and Contraction Mapping
We consider the space $\mathcal{V}$ of all bounded real-valued functions $V:\mathcal{S}\rightarrow\mathbb{R}$ defined over the state space $\mathcal{S}$. For any function $V\in\mathcal{V}$, its $L_{\infty}$-norm (or max-norm) is given by $\|V\|_{\infty}=\sup_{s\in\mathcal{S}}|V(s)|$. The space $(\mathcal{V},\|\cdot\|_{\infty})$ is a complete metric space (a Banach space).
The Bellman optimality operator $B^{*}:\mathcal{V}\rightarrow\mathcal{V}$ for the worst-path objective, with discount factor $\gamma\in(0,1)$, is defined for any value function $V\in\mathcal{V}$ and any state $s\in\mathcal{S}$ as:
$(B^{*}V)(s)=r(s)+\gamma(1-r(s))\max_{a\in\mathcal{A}}\left[\min_{s^{\prime}\in\mathcal{T}(s,a)}V(s^{\prime})\right].$ (19)
Note that if $s\in\mathcal{S}_{bb}$, then $r(s)=1$, so 1r(s)=01-r(s)=0, and $(B^{*}V)(s)=r(s)$. If $s\notin\mathcal{S}_{bb}$, then $r(s)=0$, so 1r(s)=11-r(s)=1, and $(B^{*}V)(s)=\gamma\max_{a\in\mathcal{A}}\left[\min_{s^{\prime}\in\mathcal{T}(s,a)}V(s^{\prime})\right]$. The optimal value function $V^{*}$ is the unique fixed point of this operator, i.e., $V^{*}=B^{*}V^{*}$.
Proposition 5.
The Bellman optimality operator $B^{*}$ defined in Eq. (19) is a $\gamma$-contraction mapping with respect to the $L_{\infty}$-norm. That is, for any $V_{1},V_{2}\in\mathcal{V}$:
$\|B^{*}V_{1}-B^{*}V_{2}\|_{\infty}\leq\gamma\|V_{1}-V_{2}\|_{\infty}.$
Proof.
Let $V_{1},V_{2}\in\mathcal{V}$ be two arbitrary bounded value functions. We consider any state $s\in\mathcal{S}$.
Case 1: $s\in\mathcal{S}_{bb}$. In this scenario, $(B^{*}V_{1})(s)=r(s)$ and $(B^{*}V_{2})(s)=r(s)$, as the second term in Eq. (19) (involving the maximisation) vanishes because 1r(s)=01-r(s)=0. Thus, $|(B^{*}V_{1})(s)-(B^{*}V_{2})(s)|=0$. The contraction inequality $0\leq\gamma\|V_{1}-V_{2}\|_{\infty}$ therefore holds trivially.
Case 2: $s\notin\mathcal{S}_{bb}$. In this case, $r(s)=0$, so 1r(s)=11-r(s)=1. Then,
$(B^{*}V_{1})(s)=\gamma\max_{a\in\mathcal{A}}\left[\min_{s^{\prime}\in\mathcal{T}(s,a)}V_{1}(s^{\prime})\right],$
$(B^{*}V_{2})(s)=\gamma\max_{a\in\mathcal{A}}\left[\min_{s^{\prime}\in\mathcal{T}(s,a)}V_{2}(s^{\prime})\right].$
Therefore,
$|(B^{*}V_{1})(s)-(B^{*}V_{2})(s)|=\gamma\left|\max_{a\in\mathcal{A}}\left[\min_{s^{\prime}\in\mathcal{T}(s,a)}V_{1}(s^{\prime})\right]-\max_{a\in\mathcal{A}}\left[\min_{s^{\prime}\in\mathcal{T}(s,a)}V_{2}(s^{\prime})\right]\right|.$
Let $f_{V}(a)=\min_{s^{\prime}\in\mathcal{T}(s,a)}V(s^{\prime})$. The expression is $\gamma|\max_{a}f_{V_{1}}(a)-\max_{a}f_{V_{2}}(a)|$. Using the property that for any functions $g_{1},g_{2}$, $|\max_{x}g_{1}(x)-\max_{x}g_{2}(x)|\leq\sup_{x}|g_{1}(x)-g_{2}(x)|$, we have:
$\left|\max_{a\in\mathcal{A}}f_{V_{1}}(a)-\max_{a\in\mathcal{A}}f_{V_{2}}(a)\right|\leq\sup_{a\in\mathcal{A}}|f_{V_{1}}(a)-f_{V_{2}}(a)|$
$\sup_{a\in\mathcal{A}}\left|\min_{s^{\prime}\in\mathcal{T}(s,a)}V_{1}(s^{\prime})-\min_{s^{\prime}\in\mathcal{T}(s,a)}V_{2}(s^{\prime})\right|.$
Using the property that for any functions $h_{1},h_{2}$, $|\min_{y}h_{1}(y)-\min_{y}h_{2}(y)|\leq\sup_{y}|h_{1}(y)-h_{2}(y)|$:
$\left|\min_{s^{\prime}\in\mathcal{T}(s,a)}V_{1}(s^{\prime})-\min_{s^{\prime}\in\mathcal{T}(s,a)}V_{2}(s^{\prime})\right|\leq\sup_{s^{\prime}\in\mathcal{T}(s,a)}|V_{1}(s^{\prime})-V_{2}(s^{\prime})|.$
Combining these,
$|(B^{*}V_{1})(s)-(B^{*}V_{2})(s)|$ $\leq\gamma\sup_{a\in\mathcal{A}}\left[\sup_{s^{\prime}\in\mathcal{T}(s,a)}|V_{1}(s^{\prime})-V_{2}(s^{\prime})|\right]$
$\leq\gamma\sup_{s^{\prime\prime}\in\mathcal{S}}|V_{1}(s^{\prime\prime})-V_{2}(s^{\prime\prime})|\quad(\text{as }\mathcal{T}(s,a)\subseteq\mathcal{S}\text{ and we take sup over }a)$
$=\gamma\|V_{1}-V_{2}\|_{\infty}.$
This inequality holds for any $s\notin\mathcal{S}_{bb}$.
Combining Case 1 and Case 2, for all $s\in\mathcal{S}$:
$|(B^{*}V_{1})(s)-(B^{*}V_{2})(s)|\leq\gamma\|V_{1}-V_{2}\|_{\infty}.$
Taking the supremum over all $s\in\mathcal{S}$ on the left side:
$\|B^{*}V_{1}-B^{*}V_{2}\|_{\infty}=\sup_{s\in\mathcal{S}}|(B^{*}V_{1})(s)-(B^{*}V_{2})(s)|\leq\gamma\|V_{1}-V_{2}\|_{\infty}.$
Since $0<\gamma<1$, $B^{*}$ is a $\gamma$-contraction mapping in $(\mathcal{V},\|\cdot\|_{\infty})$. The space $\mathcal{V}$ of bounded real-valued functions on $\mathcal{S}$, equipped with the $L_{\infty}$-norm, is a complete metric space (a Banach space). Therefore, by the Banach fixed-point theorem, $B^{*}$ has a unique fixed point $V^{*}\in\mathcal{V}$ such that $V^{*}=B^{*}V^{*}$. Furthermore, for any initial bounded value function $V_{0}\in\mathcal{V}$, the sequence $V_{k+1}=B^{*}V_{k}$ (i.e., value iteration) converges to $V^{*}$. The existence of a unique $V^{*}$ implies the existence of at least one stationary deterministic optimal policy $\pi^{*}$ such that for $s\notin\mathcal{S}_{bb}$:
$\pi^{*}(s)\in\arg\max_{a\in\mathcal{A}}\left[\min_{s^{\prime}\in\mathcal{T}(s,a)}V^{*}(s^{\prime})\right].$
C.4 Proof of Proposition 4
Proof.
We begin by recalling a standard identity from policy improvement theory, which relates the value difference between two policies $\pi^{i+1}$ and $\pi^{i}$ to the expected advantage under $\pi^{i+1}$:
$V^{\pi^{i+1}}(s)-V^{\pi^{i}}(s)=\mathbb{E}_{a\sim\pi^{i+1}(\cdot\mid s)}\left[A^{\pi^{i}}(s,a)\right],$
where $A^{\pi^{i}}(s,a)=Q^{\pi^{i}}(s,a)-V^{\pi^{i}}(s)$ is the advantage function under the behaviour policy $\pi^{i}$.
The update rule defined by the weighted imitation objective yields a new policy $\pi^{i+1}$ that is proportional to the exponentiated advantage:
$\pi^{i+1}(a\mid s)=\frac{\pi^{i}(a\mid s)\cdot\exp(\beta A^{\pi^{i}}(s,a))}{Z(s)},$
where $Z(s)=\sum_{a^{\prime}}\pi^{i}(a^{\prime}\mid s)\cdot\exp(\beta A^{\pi^{i}}(s,a^{\prime}))$ is a normalizing constant. This update ensures that $\pi^{i+1}$ remains in the same support set as $\pi^{i}$.
Substituting into the value difference identity, we get:
$V^{\pi^{i+1}}(s)-V^{\pi^{i}}(s)=\sum_{a}\pi^{i+1}(a\mid s)\cdot A^{\pi^{i}}(s,a)=\frac{\sum_{a}\pi^{i}(a\mid s)\cdot\exp(\beta A^{\pi^{i}}(s,a))\cdot A^{\pi^{i}}(s,a)}{\sum_{a}\pi^{i}(a\mid s)\cdot\exp(\beta A^{\pi^{i}}(s,a))}.$
This expression is a weighted average of the advantages $A^{\pi^{i}}(s,a)$, where the weights are strictly positive and increasing in the advantage. Therefore, unless all advantages are exactly zero, the expectation is strictly non-negative:
$V^{\pi^{i+1}}(s)\geq V^{\pi^{i}}(s),\quad\forall s\in\mathcal{S}.$
Appendix D Graph2Edits as Single-step Model
Graph2Edits [38] considers one-step retrosynthesis as a graph-editing game. Starting from the product molecule, the model autoregressively predicts a short sequence of primitive edits — Delete Bond, Change Bond, Change Atom, Attach Leaving Group, and finally a Terminate token. At each step the current intermediate graph is embedded by a directed message-passing neural network (D-MPNN), whose atom- and bond-level embeddings are fed to three linear heads that score every possible bond edit, atom edit and the termination symbol. The highest-scoring edit is applied to yield the next intermediate, and the process repeats until Terminate is chosen, at which point the intermediates have been fully converted to the predicted reactants. This edit vocabulary, mined from training data, covers $99.9\%$ of USPTO-50k reactions and encodes stereochemistry as well as functional leaving groups.
Trained with teacher forcing, the system already achieves a $55\%$ top-1 exact-match accuracy on USPTO-50k and maintains high validity and diversity even for long or stereochemically rich reactions; these strengths make it an ideal single-step model inside our multi-step InterRetro framework.
Appendix E Broader Impacts
By generating complete retrosynthetic routes in a single forward pass, InterRetro can shorten early-stage discovery cycles for drug discovery and materials design. The method’s tendency to propose shorter, higher-yielding sequences also points toward reduced solvent usage, energy consumption, and hazardous waste, aligning with global goals for greener and more sustainable chemistry. However, a search-free planner lowers technical barriers for malicious actors who might design routes to controlled substances or chemical weapons, and the model’s assumption of universal feasibility could mislead non-experts into attempting unsafe or impractical experiments. Therefore, we advise all users to carefully weigh the positive and negative societal impacts before employing InterRetro, and to pair any deployment with rigorous safety and feasibility checks.